# ACT-ing on inflation: Implications of non Bunch-Davies initial condition and reheating on single-field slow roll models

### v1

@@ V Slow roll models
Figure 1: The prediction of $n_{{}_{\mathrm{S}}}-r$ for $\alpha$-attractor E-model (top left) and $\alpha$-attractor T-model (top right) are plotted along with the constraints provided by P-LB-BK18 (blue) and P-ACT-LB-BK18 (yellow). The darker and light shaded areas in the contour indicate $1\sigma$ region at $68\%$ C.L. and $2\sigma$ region at $95\%$ C.L. The dashed blue, red, magenta and purple lines correspond to $\alpha=1$, $5$, $10$ and $15$ respectively. The two bounds in black correspond to the highest and lowest allowed $T_{\mathrm{re}}$. The gray and green shaded areas correspond to $n=2$ and $n=8$. The middle plots correspond to the value of $\beta_{0}$ for $\alpha$-attractor E-model (middle left) and for $\alpha$-attractor T-model (middle right), that is required in order to satisfy the mean value of $n_{{}_{\mathrm{S}}}=0.9743$ as provided by the current ACT data. In the bottom plots, the tensor-to-scalar ratios are plotted which are obtained after incorporating the modification from the NBD initial state in order to satisfy $n_{{}_{\mathrm{S}}}=0.9743$.
@@ V.1 $\alpha$-attractor models
×$\quad\times\left[\exp\left(\sqrt{\frac{2}{3\alpha}}\frac{\phi}{M_{{}_{\mathrm{Pl}}}}\right)-1\right]^{-2}$ (26)
@@
×$\quad\times\left[(n-1)\frac{\mathrm{sech}^{2}\left({\frac{1}{\sqrt{6\alpha}}}\frac{\phi}{M_{{}_{\mathrm{Pl}}}}\right)}{\mathrm{tanh}^{2}\left({\frac{1}{\sqrt{6\alpha}}}\frac{\phi}{M_{{}_{\mathrm{Pl}}}}\right)}-2\right].$ (28)
@@
A recent study has demonstrated that, in light of the P-ACT-LB-BK18 data, the Starobinsky model of inflation is pushed toward the edge of the allowed $n_{{}_{\mathrm{S}}}-r$ region, lying close to the $2\sigma$ contour. We begin by investigating whether this tension can be alleviated through a subsequent phase of reheating. At this point, it is important to highlight a key consideration: in such analyses, the reheating EoS parameter can often be treated as a free parameter. We will adopt this approach for most of the models considered in this work. However, in the case of the $\alpha$-attractor models, for simplicity and to reduce the number of free parameters, we shall assume that the reheating EoS is determined by the behavior of the inflaton field near the minimum of the potential. Specifically, if the potential near the minimum behaves as $V(\phi)\sim\phi^{n}$, then the effective EoS is given by $w=(n-2)/(n+2)$ [73].
@@
Figure 2: The prediction of $n_{{}_{\mathrm{S}}}-r$ for $\phi^{n}$ potential (top left) and hilltop potential (top right) are plotted along with the constraints provided by P-LB-BK18 (blue) and P-ACT-LB-BK18 (yellow). The darker and light shaded areas in the contour indicate $1\sigma$ region at $68\%$ C.L. and $2\sigma$ region at $95\%$ C.L. The dashed blue, red, magenta lines correspond to $n=1/2$, $1/3$, and $2/3$ in $\phi^{n}$ potential and the dashed blue, red, magenta, purple lines correspond to $f/M_{{}_{\mathrm{Pl}}}=10$, $12$, $15$ and $20$ in hilltop model. The two bounds in black correspond to the highest and lowest allowed $T_{\mathrm{re}}$. The gray and green shaded areas correspond to $w=0$ and $w=0.9$. The middle plots correspond to the value of $\beta_{0}$ for $\phi^{n}$ potential (middle left) and for hilltop potential (middle right), that is required in order to satisfy the mean value of $n_{{}_{\mathrm{S}}}=0.9743$ as provided by the current ACT data. In the bottom plots, the tensor-to-scalar ratios are plotted which are obtained after incorporating the modification from the NBD initial state in order to satisfy $n_{{}_{\mathrm{S}}}=0.9743$.
@@ V.2 $\phi^{n}$ potential
In the top left panel of Fig. 2, we show the predictions for the scalar spectral index $n_{{}_{\mathrm{S}}}$ and tensor-to-scalar ratio $r$ in chaotic inflation models for two representative values of the reheating EoS parameter: $w=0$ and $w=0.9$. The observational contours from P-LB-BK18 and P-ACT-LB-BK18 are also displayed for comparison. As noted earlier, the black curves delineate the parameter space bounded by the maximum and minimum allowed reheating temperatures, $T_{\mathrm{re}}$. Previously, in the absence of ACT data, chaotic inflation was nearly ruled out, even for small values of $n$. However, the inclusion of ACT, which shifts the central value of $n_{{}_{\mathrm{S}}}$ upward, has significantly altered the viability of these models. We find that for $n=1/5$ and even $1/3$, the models now fall within the $1\sigma$ confidence region for lower values of $w$. The case with $n=2/3$ lies on the edge of the $2\sigma$ contour, while the conventional $\phi^{2}$ model remains well outside the allowed region. Interestingly, recent studies have demonstrated that introducing a non-minimal coupling between the inflaton and gravity can shift the predictions of the $\phi^{2}$ model into the allowed observational region [12].
We now investigate whether an NBD initial state can help achieve the target spectral index $n_{{}_{\mathrm{S}}}=0.9743$ as indicated by the P-ACT-LB-BK18 data. To this end, we compute the required values of the NBD parameter $\beta_{0}$, which are presented in the middle left panel of Fig. 2 as a function of the potential index $n$. In some cases, achieving the desired value of $n_{{}_{\mathrm{S}}}$ requires lowering it from the value obtained under the standard BD vacuum, which in turn leads to negative values of $\beta_{0}$. The corresponding predictions for the tensor-to-scalar ratio $r$, including the NBD corrections, are shown in the bottom left panel of Fig. 2. We observe that $r$ tends to exceed the current upper bounds in this scenario. Notably, even the case with $n=1/3$, which initially fell within the $1\sigma$ region under BD initial conditions, now predicts a larger $r$ due to the downward adjustment of $n_{{}_{\mathrm{S}}}$, thereby moving outside the allowed region. Therefore, for monomial potentials, the introduction of NBD initial conditions does not significantly improve consistency with the $n_{{}_{\mathrm{S}}}-r$ observational bounds. In such cases, it may be more promising to explore models with non-canonical kinetic terms.
@@ V.3 Hilltop inflation
Hilltop inflation refers to a class of models where the inflaton field starts near the top of a potential hill and rolls slowly towards the minimum, enabling a prolonged period of inflation. These models are characterized by potentials with a maximum at or near the origin, typically of the form, $V=V0[1-(\phi/f)^{n}+\cdots]$ where $f$ sets the width of the hill and $n$ determines the steepness. Due to the flatness near the top, hilltop models can naturally produce a low tensor-to-scalar ratio $r$, often within observational limits. In this work we shall consider the potential to be [76, 77, 78, 79]
@@
In the top right panel of Fig. 2, we present the predictions for $n_{{}_{\mathrm{S}}}$$r$ in hilltop inflation for two representative values of the reheating EoS parameter: $w=0$ and $w=0.9$. For reference, we overlay the observational contours from P-LB-BK18 and P-ACT-LB-BK18. As in previous cases, the black curves show the boundaries of the parameter space set by the maximum and minimum allowed reheating temperatures, $T_{\mathrm{re}}$. We find that the hilltop potential remains compatible with the $n_{{}_{\mathrm{S}}}-r$ bounds, even after the ACT data shift the preferred value of $n_{{}_{\mathrm{S}}}$ to higher values. To meet this updated constraint, a larger value of the parameter $f$ is needed. From the figure, we observe that the model satisfies the $1\sigma$ bound for values of $f$ in the range $12M_{{}_{\mathrm{Pl}}}\lesssim f\lesssim 15M_{{}_{\mathrm{Pl}}}$.
Although hilltop inflation can satisfy the $n_{{}_{\mathrm{S}}}$$r$ constraint even with the standard BD initial state, we examine the impact of a NBD initial condition for completeness. In the middle panel of Fig. 2, we show the values of the $\beta_{0}$ required to achieve $n_{{}_{\mathrm{S}}}=0.9743$. We observe that for a steep reheating EoS parameter $w=0.9$ and for $f\gtrsim 17M_{{}_{\mathrm{Pl}}}$, the value of $\beta_{0}$ becomes negative in order to lower $n_{{}_{\mathrm{S}}}$ from a higher value predicted under the BD assumption. The corresponding values of $r$, incorporating the effects of the NBD correction, are displayed in the bottom right panel of Fig. 2. It is evident that the inclusion of the NBD initial state allows a broader range of the parameter $f$ to remain consistent with observational bounds on $r$.
@@ V.4 Natural inflation
It is well known that natural inflation is under increasing tension with current cosmological observations, particularly due to its prediction of a relatively large $r$, which is disfavored by the latest data from P-ACT-LB-BK18. Despite its appealing theoretical motivations, the model struggles to remain within the favored region of the $n_{{}_{\mathrm{S}}}-r$ parameter space [82]. In this work, we investigate whether introducing a NBD initial state can help reconcile natural inflation with current observational bounds. Specifically, we aim to achieve the desired spectral index $n_{{}_{\mathrm{S}}}=0.9743$, as indicated by the P-ACT-LB-BK18 data, by tuning the parameter $\beta_{0}$. Our analysis reveals that in order to match the observed value of $n_{{}_{\mathrm{S}}}$, a relatively large value of $\beta\geq 0.6$ is required. However, this modification comes at a cost. The introduction of the NBD correction significantly enhances the predicted value of the $r$, pushing it above the observational upper limit of $r\sim 0.05$. Thus, even with the inclusion of a non-standard initial state, natural inflation remains in tension with the current data, indicating that the NBD initial condition does not resolve the fundamental issue.
VI Conclusions

### v2

@@ V Slow roll models
Figure 1: The prediction of $n_{{}_{\mathrm{S}}}-r$ for $\alpha$-attractor E-model (top left) and $\alpha$-attractor T-model (top right) are plotted along with the constraints provided by P-LB-BK18 (blue) and P-ACT-LB-BK18 (yellow). The darker and light shaded areas in the contour indicate $1\sigma$ region at $68\%$ C.L. and $2\sigma$ region at $95\%$ C.L. Note that the constraints are drawn using the data of [8] as also given in [21]. The dashed blue, red, magenta and purple lines correspond to $\alpha=1$, $5$, 1010 and 1515 respectively. The two bounds in black correspond to the highest and lowest allowed $T_{\mathrm{re}}$. The gray and green shaded areas correspond to $n=2$ and $n=8$. The middle plots correspond to the value of $\beta_{0}$ for $\alpha$-attractor E-model (middle left) and for $\alpha$-attractor T-model (middle right), that is required in order to satisfy the mean value of $n_{{}_{\mathrm{S}}}=0.9743$ as provided by the current ACT data. In the bottom plots, the tensor-to-scalar ratios are plotted which are obtained after incorporating the modification from the NBD initial state in order to satisfy $n_{{}_{\mathrm{S}}}=0.9743$.
@@ V.1 $\alpha$-attractor models
$\quad\times\left[\exp\left(\sqrt{\frac{2}{3\alpha}}\frac{\phi}{M_{{}_{\mathrm{Pl}}}}\right)-1\right]^{-2}$ (26)
@@
$\quad\times\left[(n-1)\frac{\mathrm{sech}^{2}\left({\frac{1}{\sqrt{6\alpha}}}\frac{\phi}{M_{{}_{\mathrm{Pl}}}}\right)}{\mathrm{tanh}^{2}\left({\frac{1}{\sqrt{6\alpha}}}\frac{\phi}{M_{{}_{\mathrm{Pl}}}}\right)}-2\right].$ (28)
@@
A recent study has demonstrated that, in light of the P-ACT-LB-BK18 data, the Starobinsky model of inflation [20] is pushed toward the edge of the allowed $n_{{}_{\mathrm{S}}}-r$ region, lying close to the $2\sigma$ contour [21]. We begin by investigating whether this tension can be alleviated through a subsequent phase of reheating. At this point, it is important to highlight a key consideration: in such analyses, the reheating EoS parameter can often be treated as a free parameter, without directly linking $w_{\mathrm{re}}=(n-2)/(n+2)$ for a potential behaving $\phi^{n}$ at minima. In such cases it is considered that after the end of inflation, the potential has different behavior at the minima. We will adopt this approach for most of the models considered in this work. However, in the case of the $\alpha$-attractor models, for simplicity and to reduce the number of free parameters, we shall assume that the reheating EoS is determined by the behavior of the inflaton field near the minimum of the potential. Specifically, if the potential near the minimum behaves as $V(\phi)\sim\phi^{n}$, then the effective EoS is given by $w_{\mathrm{re}}=(n-2)/(n+2)$ [77].
@@
Figure 2: The prediction of $n_{{}_{\mathrm{S}}}-r$ for $\phi^{n}$ potential (top left) and hilltop potential (top right) are plotted along with the constraints provided by P-LB-BK18 (blue) and P-ACT-LB-BK18 (yellow). The darker and light shaded areas in the contour indicate $1\sigma$ region at $68\%$ C.L. and $2\sigma$ region at $95\%$ C.L. The dashed blue, red, magenta lines correspond to $n=1/2$, $1/3$, and $2/3$ in $\phi^{n}$ potential and the dashed blue, red, magenta, purple lines correspond to $f/M_{{}_{\mathrm{Pl}}}=10$, $12$, 1515 and $20$ in hilltop model. The two bounds in black correspond to the highest and lowest allowed $T_{\mathrm{re}}$. The gray and green shaded areas correspond to $w_{\mathrm{re}}=0$ and $w_{\mathrm{re}}=0.9$. The middle plots correspond to the value of $\beta_{0}$ for $\phi^{n}$ potential (middle left) and for hilltop potential (middle right), that is required in order to satisfy the mean value of $n_{{}_{\mathrm{S}}}=0.9743$ as provided by the current ACT data. In the bottom plots, the tensor-to-scalar ratios are plotted which are obtained after incorporating the modification from the NBD initial state in order to satisfy $n_{{}_{\mathrm{S}}}=0.9743$.
@@ V.2 $\phi^{n}$ potential
In the top left panel of Fig. 2, we show the predictions for the scalar spectral index $n_{{}_{\mathrm{S}}}$ and tensor-to-scalar ratio $r$ in chaotic inflation models for two representative values of the reheating EoS parameter: $w_{\mathrm{re}}=0$ and $w_{\mathrm{re}}=0.9$. The observational contours from P-LB-BK18 and P-ACT-LB-BK18 are also displayed for comparison. As noted earlier, the black curves delineate the parameter space bounded by the maximum and minimum allowed reheating temperatures, $T_{\mathrm{re}}$. Previously, in the absence of ACT data, chaotic inflation was nearly ruled out, even for small values of $n$ [5, 6, 12, 13]. However, the inclusion of ACT, which shifts the central value of $n_{{}_{\mathrm{S}}}$ upward, has significantly altered the viability of these models. We find that for $n=1/5$ and even $1/3$, the models now fall within the $1\sigma$ confidence region for lower values of $w_{\mathrm{re}}$. The case with $n=2/3$ lies on the edge of the $2\sigma$ contour, while the conventional $\phi^{2}$ model remains well outside the allowed region. Interestingly, recent studies have demonstrated that introducing a non-minimal coupling between the inflaton and gravity can shift the predictions of the $\phi^{2}$ model into the allowed observational region [21].
We now investigate whether an NBD initial state can help achieve the target spectral index $n_{{}_{\mathrm{S}}}=0.9743$ as indicated by the P-ACT-LB-BK18 data. To this end, we compute the required values of the NBD parameter $\beta_{0}$, which are presented in the middle left panel of Fig. 2 as a function of the potential index $n$. In some cases, achieving the desired value of $n_{{}_{\mathrm{S}}}$ requires lowering it from the value obtained under the standard BD vacuum, which in turn leads to negative values of $\beta_{0}$. The corresponding predictions for the tensor-to-scalar ratio $r$, including the NBD corrections, are shown in the bottom left panel of Fig. 2. We observe that $r$ tends to exceed the current upper bounds in this scenario. Notably, even the case with $n=1/3$, which initially fell within the $1\sigma$ region under BD initial conditions, now predicts a larger $r$ due to the downward adjustment of $n_{{}_{\mathrm{S}}}$, thereby moving outside the allowed region. Therefore, for monomial potentials, the introduction of NBD initial conditions does not significantly improve consistency with the $n_{{}_{\mathrm{S}}}-r$ observational bounds. In such cases, it may be more promising to explore models with non-canonical kinetic terms [99, 100, 101, 21].
@@ V.3 Hilltop inflation
Hilltop inflation refers to a class of models where the inflaton field starts near the top of a potential hill and rolls slowly towards the minimum, enabling a prolonged period of inflation. These models are characterized by potentials with a maximum at or near the origin, typically of the form, $V=V0[1-(\phi/f)^{n}+\cdots]$ where $f$ sets the width of the hill and $n$ determines the steepness. Due to the flatness near the top, hilltop models can naturally produce a low tensor-to-scalar ratio $r$, often within observational limits. In this work we shall consider the potential to be [102, 103, 104, 105, 106, 107, 108]
@@
In the top right panel of Fig. 2, we present the predictions for $n_{{}_{\mathrm{S}}}$$r$ in hilltop inflation for two representative values of the reheating EoS parameter: $w_{\mathrm{re}}=0$ and $w_{\mathrm{re}}=0.9$. For reference, we overlay the observational contours from P-LB-BK18 and P-ACT-LB-BK18. As in previous cases, the black curves show the boundaries of the parameter space set by the maximum and minimum allowed reheating temperatures, $T_{\mathrm{re}}$. We find that the hilltop potential remains compatible with the $n_{{}_{\mathrm{S}}}-r$ bounds, even after the ACT data shift the preferred value of $n_{{}_{\mathrm{S}}}$ to higher values. To meet this updated constraint, a larger value of the parameter $f$ is needed. From the figure, we observe that the model satisfies the $1\sigma$ bound for values of $f$ in the range $12M_{{}_{\mathrm{Pl}}}\lesssim f\lesssim 15M_{{}_{\mathrm{Pl}}}$.
Although hilltop inflation can satisfy the $n_{{}_{\mathrm{S}}}$$r$ constraint even with the standard BD initial state, we examine the impact of a NBD initial condition for completeness. In the middle panel of Fig. 2, we show the values of the $\beta_{0}$ required to achieve $n_{{}_{\mathrm{S}}}=0.9743$. We observe that for a steep reheating EoS parameter $w_{\mathrm{re}}=0.9$ and for $f\gtrsim 17M_{{}_{\mathrm{Pl}}}$, the value of $\beta_{0}$ becomes negative in order to lower $n_{{}_{\mathrm{S}}}$ from a higher value predicted under the BD assumption. The corresponding values of $r$, incorporating the effects of the NBD correction, are displayed in the bottom right panel of Fig. 2. It is evident that the inclusion of the NBD initial state allows a broader range of the parameter $f$ to remain consistent with observational bounds on $r$.
@@ V.4 Natural inflation
It is well known that natural inflation [112] is under increasing tension with current cosmological observations, particularly due to its prediction of a relatively large $r$ [113, 114, 115], which is disfavored by the latest data from P-ACT-LB-BK18. Despite its appealing theoretical motivations [^2] The flatness of the potential in natural inflation is protected by the shift symmetry, which naturally suppresses large radiative corrections, solving the $\eta$-problem. It is motivated by axion physics, where many string theory compactifications produce axion-like fields. Also, such a cosine-type potential in Eq. (37) naturally arises in theories with spontaneous breaking of a global symmetry. The potential also sustain inflation without excessive fine-tuning. For detailed discussion regarding this, see [116, 117, 118, 111]., the model struggles to remain within the favored region of the $n_{{}_{\mathrm{S}}}-r$ parameter space [111]. In this work, we investigate whether introducing a NBD initial state can help reconcile natural inflation with current observational bounds. Specifically, we aim to achieve the desired spectral index $n_{{}_{\mathrm{S}}}=0.9743$, as indicated by the P-ACT-LB-BK18 data, by tuning the parameter $\beta_{0}$. Our analysis reveals that in order to match the observed value of $n_{{}_{\mathrm{S}}}$, a relatively large value of $\beta\geq 0.6$ is required. However, this modification comes at a cost. The introduction of the NBD correction significantly enhances the predicted value of the $r$, pushing it above the observational upper limit of $r\sim 0.05$. Thus, even with the inclusion of a non-standard initial state, natural inflation remains in tension with the current data, indicating that the NBD initial condition does not resolve the fundamental issue.
VI How large $\beta_{k}$ can be?
In this section we shall discuss the viability of choosing the value of the NBD parameter $\beta_{k}$ to avoid backreaction. As already mentioned, from the gravitational backreaction, we have an upper bound on $\beta_{k}$ to be $\beta_{k}\lesssim\sqrt{\epsilon\eta^{\prime}}HM_{{}_{\mathrm{Pl}}}/M^{2}$, where $M$ defines the scale of new physics [74, 75, 76]. From the choice of our $\beta_{k}$ in Eq. (11), it is evident that the change in the power spectrum is controlled by $k/(aM)\sim H/M$. Hence maximum allowed value of $\beta_{0}$ can be controlled by changing the value of $H/M$. For consistency with effective field theory, $M>H$ is typically required. In this work we have chosen at pivot scale, $H_{\ast}/M=0.2$ as suggested in [42]. For a typical choice of $\epsilon\sim\eta^{\prime}\sim 10^{-2}$ and $H\sim 10^{-5}M_{{}_{\mathrm{Pl}}}$ the allowed value of $\beta_{k}$ can even reach of the order of $\mathcal{O}(10)$. Which can also be understood from figure 1 of [75]. In such a case, our choice of $\beta_{0}<0.2$ lies well below the allowed value.
There is yet another issue associated with the largeness of $\beta_{k}$ that we shall discuss. This comes from the amount of non-Gaussianity that is produced due to the NBD initial state [47, 68, 119, 69, 120]. The non-Gaussianity parameter associated to three scalar mode is given by the ratio to the three-point correlation to the power spectrum as
$f_{{}_{\mathrm{NL}}}$ $=-\frac{10}{3}\frac{1}{(2\pi)^{4}}{G_{\zeta\zeta\zeta}(\bm{k}_{1},\bm{k}_{2},\bm{k}_{3})}$ (41)
$\quad\times\left[\frac{\mathcal{P}_{{}_{\mathrm{S}}}(k_{1})}{k_{1}^{3}}\frac{\mathcal{P}_{{}_{\mathrm{S}}}(k_{2})}{k_{2}^{3}}+\text{two combinations}\right]^{-1},$ (42)
where $G_{\zeta\zeta\zeta}(\bm{k}_{1},\bm{k}_{2},\bm{k}_{3})$ is the three-point correlation of the perturbation $\hat{\zeta}$, defined as
$\langle\hat{\zeta}_{\bm{k}_{1}}\hat{\zeta}_{\bm{k}_{2}}\hat{\zeta}_{\bm{k}_{3}}\rangle=\frac{1}{(2\pi)^{3/2}}G_{\zeta\zeta\zeta}(\bm{k}_{1},\bm{k}_{2},\bm{k}_{3})\delta^{(3)}(\bm{k}_{1}+\bm{k}_{2}+\bm{k}_{3}).$ (43)
The observational constraints on $f_{{}_{\mathrm{NL}}}$ in local, equilateral and orthogonal shape is given by [121]
$f_{{}_{\mathrm{NL}}}^{\rm local}$ $=-0.9\pm 5.1,$ (44a)
$f_{{}_{\mathrm{NL}}}^{\rm equil}$ $=-26\pm 47,$ (44b)
$f_{{}_{\mathrm{NL}}}^{\rm ortho}$ $=-38\pm 24.$ (44c)
It is well known that a change in $\beta_{k}$ can significantly affect $f_{{}_{\mathrm{NL}}}$. Hence it becomes important to check that our choice of $\beta_{k}$ does not produce an $f_{{}_{\mathrm{NL}}}$ beyond the allowed range. The standard consistency condition does not hold if the initial state is not BD. The NBD state can produce non-Gaussianity which can be dominant in the flattened shape, where $k_{1}=k_{2}=k_{3}/2$ [68]. The non-Gaussianity parameter in flattened shape is given by
$f_{{}_{\mathrm{NL}}}^{\rm flattened}\sim\frac{\dot{\phi}^{2}}{M^{4}}\beta_{k},$ (45)
where after taking into account the bound from gravitational backreaction, the maximum expected value of $f_{{}_{\mathrm{NL}}}$ reads
$f_{{}_{\mathrm{NL}}}^{\rm flattened}\sim\epsilon\sqrt{\epsilon\eta^{\prime}}\left(\frac{H}{M_{{}_{\mathrm{Pl}}}}\right)^{2}\left(\frac{M_{{}_{\mathrm{Pl}}}}{M}\right)^{5}.$ (46)
Now, taking the typical value of the parameters $\epsilon\sim\eta\sim 10^{-2}$ and $H\sim 10^{-5}M_{{}_{\mathrm{Pl}}}$ we see that, the choice of $M=5H$ at pivot scale, produces large $f_{{}_{\mathrm{NL}}}$. One can reduce $f_{{}_{\mathrm{NL}}}$ down to order $\mathcal{O}(10)$, by choosing an lower value of $H/M$, say for instance, taking $M\sim 70H$. However, adopting such a value of $M$, can exclude parts of the parameter space in $w_{\mathrm{re}}-T_{\mathrm{re}}$, which might otherwise be viable for lower values of $\epsilon$, $\eta^{\prime}$. A deeper analysis of specific inflationary models is required to assess these possibilities. Hence, it is safe to say that, for some cases, $n_{{}_{\mathrm{S}}}$ can not be brought back to the allowed $1\sigma$ region solely by invoking a NBD initial state, without violating the observational bound on $r$, or generating an unacceptably large $f_{{}_{\mathrm{NL}}}$. As an alternative NBD state, one may further explore another class of states called coherent states which possibly has lesser issues due to large non-Gaussianities and backreaction [71, 122, 123, 124].
VII Conclusions